Find the equation of the ellipse in standard form whose Foci is (2,7) and (-2,7) and the length of the major axis is 6


1
Expert's answer
2021-10-20T00:10:02-0400

The center is located midway between the vertices on the major axis:

center="(\frac{(2+(-2)}{2},\frac{7+7}{2})=(0,7)"

length of major axis= 6=2a

"\implies a=3"

"c^2=a^2-b^2\\
c=2,a=3\\
2^2=3^2-b^2\\
b^2=9-4=5\\
b=\sqrt 5"

The standard form equation of the ellipse is "\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1"

"\frac{x^2}{9}+\frac{(y-7)^2}{5}=1"